OpenStudy (anonymous):

What are the real or imaginary solutions of 125x^3+343=0?

OpenStudy (anonymous):

\[125x ^{3}+343=0\]

OpenStudy (dangerousjesse):

Real solution:

OpenStudy (dangerousjesse):

Solve for x over the real numbers:\[125 x^3+343 = 0 \]Isolate terms with x to the left hand side. Subtract 343 from both sides:\[125 x^3 = -343 \]Divide both sides by a constant to simplify the equation. Divide both sides by 125:\[x^3 = \frac{-343}{125 }\]Elminate the power on the left hand side. Take cube roots of both sides: What's your answer?

OpenStudy (anonymous):

\[x=\sqrt[3]{-343/125}\]

OpenStudy (anonymous):

That's probably wrong

OpenStudy (dangerousjesse):

Can you simplify that any more?

OpenStudy (anonymous):

\[x=\sqrt[3]{2.744}\]

OpenStudy (anonymous):

I'm not the best with simplifying radicals

OpenStudy (dangerousjesse):

Not quite, how about you just simplify the fraction?

OpenStudy (anonymous):

Sorry, I still don't see, -2 93/125?

OpenStudy (dangerousjesse):

Actually, the answer would just be \[-\frac{7}{5}\]

OpenStudy (anonymous):

Oh

OpenStudy (dangerousjesse):

You were supposed to completely ignore x^3 and find the cube roots of the numerator and denominator.

OpenStudy (anonymous):

7^3/5^3

OpenStudy (anonymous):

yeah I get it now the cuberoot goes to simplify the fration by getting the cubroot of the numerator and denominator.

OpenStudy (anonymous):

*fraction

OpenStudy (dangerousjesse):

Yep, you can just toss out the powers once you get to that point. Nice job.

OpenStudy (anonymous):

OK, thanks once again :D

OpenStudy (dangerousjesse):

Haha no problem :)